# Biomedical Event Extraction via Structure-aware Generation

### v1

@@ Abstract
Biomedical Event Extraction (BEE) is a critical task that involves modeling complex relationships between fine-grained entities in biomedical text data. However, most existing BEE models rely on classification methods that neglect the label semantics and argument dependency structure within the data. To address these limitations, we propose GenBEE, a generative model enhanced with a structure-aware prefix for biomedical event extraction. GenBEE constructs event prompts that leverage knowledge distilled from large language models (LLMs), thereby incorporating both label semantics and argument dependency relationships. Additionally, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes with structural prompts, enriching the generation process with structural features. Extensive experiments on three benchmark datasets demonstrate the effectiveness of GenBEE and it achieves state-of-the-art performance on the MLEE and GE11 datasets. Furthermore, our analysis shows that the structural prefixes effectively bridge the gap between structural prompts and the representation space of generative models, enabling better integration of event structural information.
@@ I Introduction
Biomedical Event Extraction (BEE) is a challenging task that involves identifying molecular events from natural language text, where an event typically includes certain biomedical entities, such as genes, proteins, or cellular components [1]. BEE is typically divided into two subtasks: (1) Biomedical Event Trigger Detection, which aims to identify trigger words and classify them by event types that represent the presence of biomedical events. (2) Biomedical Event Argument Extraction, which aims to identify the arguments and associate them with specific roles in biomedical events, correlating them with the corresponding event triggers. Moreover, BEE can provide invaluable information to facilitate the curation of knowledge bases and biomolecular pathways [2].
@@
Traditionally, BEE has been formulated as a classification problem [3, 4]. However, these classification-based methods often struggle with encoding label semantics and other forms of weak supervision information [5, 6, 7]. Moreover, many current methods extract events in a pipeline manner, where trigger detection and argument extraction are treated as separate phases, which hinders the learning of shared knowledge between subtasks [8]. Recently, generation-based models have been proposed for Event Extraction (EE) on general domains [9, 10]. These models typically cast the event extraction task as a sequence generation problem by utilizing encoder-decoder-based Pre-trained Language Models (PLMs) [11, 12] to output conditional generation sequences. These generation-based models benefit from natural language prompts to learn additional semantic information [13]. Apart from employing natural language prompts to enhance semantic learning from the given context, some recent works, such as GTEE [14] and AMPERE [15], have also proposed injecting continuous prompts [16, 17] into their frameworks for better contextual representation learning for event argument extraction.
@@
Figure 1: An illustration of the nested events identified from the given text. Nested events constitute approximately 25% of all events in biomedical text data.
@@
However, most current generation-based EE models [14, 13, 15] have overlooked events with complex structures, such as nested events. Compared to flat event structures, texts in the biomedical domain often contain more nested event structures [18]. For example, in benchmark datasets like MLEE [19] and GE11 [20], the nested event structures constitute approximately 25% and 28% of all events in the datasets, respectively. Figure 1 provides an example of two nested events identified from the given text. The trigger word “promoters” of the Positive Regulation event (E2) serves as the argument role Site for the trigger word “bind”, which triggers the Binding event (E1). Furthermore, “TCF-1 alpha” serves as the argument role Theme for both events (E1 and E2). Moreover, a trigger word may also trigger different events in a sentence, depending on its event type or arguments. Events with such structures are called overlapping events. The rest of the events are regarded as flat events. Identifying these complex structures is non-trivial, as it can reveal the underlying relationships between events. For example, a “Site2” argument in a “Binding” event is likely to be a trigger for a “Positive Regulation” event.
@@
An intuitive way to incorporate structural information through prompt engineering [21] is by adding natural language prompts that describe complex structural information as part of the input to the generative PLM’s encoder. This helps elicit the rich knowledge distributed in the PLM. However, [22] pointed out that the mainstay contextual representations in a generative language model’s encoder tend to weaken other high-level information such as structural information, thereby reducing the effects of such information at the decoder. Moreover, it is also unclear whether generative PLMs, such as BART, contain sufficient domain knowledge to represent complex biomedical events. Therefore, how to effectively incorporate structural information into generative BEE models remains an open question.
@@
To address the above challenges, we propose GenBEE (Generative Biomedical Event Extraction model with structure-aware prefix), a generative framework to fully exploit complex event structures in biomedical texts. GenBEE first constructs type description prompts to integrate the semantics of type labels. It then distills knowledge from large language models (LLMs), such as GPT-4 [23], to create event template prompts, incorporating additional weak supervision information, including dependencies among argument roles, to enhance contextual representation learning. Furthermore, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes, enriched with structural prompts, to infuse the generation process with structural features. Altogether, GenBEE serves as an end-to-end BEE framework that simultaneously tackles event trigger detection and event argument extraction, effectively leveraging the shared knowledge and dependencies between these two subtasks. The main contributions of this paper are summarized as follows:
@@
  • We propose GenBEE, a structure-aware generation model for biomedical event extraction. Our end-to-end model can distill knowledge from the LLM to incorporate label semantics, dependency relationships among arguments, and shared knowledge between subtasks.
@@
  • We introduce a structural prefix learning module to guide the generation process. This module constructs the structure-aware prefix to link continuous prompts with the representation space of generative models. Notably, this plug-and-play module can be added to any biomedical event extraction framework.
  • We evaluate GenBEE on three widely-used BEE benchmark datasets, namely MLEE, GE11, and PHEE. Extensive experimental results and analyses demonstrate the effectiveness of the proposed GenBEE model.
@@ II-A Biomedical Event Extraction
Many current methods for biomedical event extraction adopt a “pipeline” strategy, where trigger detection and argument extraction are treated as separate phases. However, “pipeline” methods may result in the propagation of errors. To address this issue, Trieu et al. [8] introduced an end-to-end method called DeepEventMine, which simultaneously identifies triggers and allocates roles to entities, thereby mitigating error propagation. However, DeepEventMine requires the full annotation of all entities, which may not always be accessible in datasets. Ramponi et al. [6] considered biomedical event extraction as a sequence labeling task that involves joint training models for trigger detection and argument extraction via multi-task learning. Aside from these classification-based methods, recent works have formulated event extraction as a question-answering task [24, 25]. This new paradigm transforms traditional classification methods into multiple questioning rounds, producing a natural language answer about a trigger or an argument in each round. Current QA-based event extraction methods are primarily focused on formulating separate questions for different events and argument types. However, these QA methods lack the dependency information between arguments, which could be helpful for assigning argument types.
@@ II-B Generative Event Extraction
Recently, another line of work that treats event extraction as a conditional generation problem [26, 27, 28, 13] has evolved. The development of generation-based event extraction models begins with the reformulation of event extraction problems as a generation task. For example, TANL [26] treated event extraction as translation tasks involving augmented natural languages, used a variety of brackets and vertical bar symbols to represent embedded language labels, and tuned the model to predict these augmented natural language labels. TempGen [27] proposed using templates for constructing role-filler tasks for entity extraction, generating outputs that fill role entities into non-natural templated sequences. To adapt generation-based models to the task of event argument extraction, BART-Gen [28] investigated generative document-level event argument extraction. However, these methods only use naive natural language prompts and do not consider incorporating semantic information into the natural language prompts.
To better incorporate label semantics into the model, DEGREE [13] proposed a template-based model featuring label semantics with designed prompts to extract event information. More recently, researchers have started to advance this line of work. For example, GTEE [14] used prefixes [16], which are sequences of tunable continuous prompts, to incorporate contextual type information into the generation-based model. AMPERE [15] proposed to equip the generative model with AMR-aware prefixes to embed the AMR graph, providing more contextual information. However, not much work investigates the effectiveness of incorporating structural features of complex events into the generative models, even though this information has proven beneficial in some extractive models. Therefore, in this paper, we investigate methods to integrate complex event structural information into generative models.
@@ III Method
In this section, we introduce our GenBEE model in detail. The proposed GenBEE model performs generative biomedical event extraction as follows: The extraction process is divided into a number of subtasks based on each pre-defined event type of a dataset. Given a Dataset $\mathcal{D}$ $=$ $\left\{\mathcal{C}_{j}\mid j\in[1,|\mathcal{D}|]\right\}$ with its pre-defined event type set $\mathcal{E}$ $=$ $\left\{e_{i}\mid i\in[1,|\mathcal{E}|]\right\}$, where $\mathcal{C}_{j}$ denotes a textual context and $e_{i}$ denotes an event type. Each subtask is denoted as $S_{e_{i},\mathcal{C}}$, for the type of event $e_{i}$ and the context $\mathcal{C}$, where $\mathcal{C}$ contains all the textual contexts in the dataset. As such, we have a total of $|\mathcal{E}|$ generation subtasks based on each event type for all textual contexts $\mathcal{C}$. The proposed model conducts training iteratively based on each subtask with all textual contexts in the dataset. Figure 2 shows the architecture of the proposed model, which consists of the following modules: Event Prompt Construction and Concatenation, Structural Prompt Construction, Prefix Generation, BART Encoding-Decoding, and Event Generation.
@@ III-A Event Prompt Construction
Most previous works [14, 15] proposed to use manually designed templates for event extraction. However, it is challenging and time-consuming to define event templates for each event type from the biomedical event datasets, which requires domain knowledge. In this work, we propose to employ the Large Language Model (LLM) GPT-4 [23] with designed prompts to generate the event templates for the event types pre-defined in the biomedical event datasets. Figure 3 shows an input prompt that we have designed for generating the event template for the “Binding” event using GPT-4. The input prompt consists of an “Instruction” that specifies the generation objective for the biomedical event extraction template and a “Basic Template” that outlines the target trigger and argument roles, constructed based on the pre-defined roles given in the technical report associated with the biomedical event dataset. The output of GPT-4 is an updated event template for the corresponding event type, which contains richer information than the basic template for generative event extraction.
@@
where “;” denotes the sequence concatenation, and $[SEP]$ is the special separator token in the PLM. For each subtask based on $e_{i}$, we construct the input tokens for all the contexts in the dataset.
@@ III-B Structural Prefix Learning
In biomedical texts, events are often complex and interrelated. Identifying the latent relationships among entities across different events could provide important clues for capturing complex event structures, thereby enhancing representation learning. Table II shows the structural prompts [29] constructed in the proposed model to capture different event structures for the Binding event. The structural prompts aim to provide descriptions of different event structures and then elicit the PLM to identify complex event structures by capturing the latent relationships among entities across different biomedical events. As shown in Table II, the structural prompts are the same for all event types, except the corresponding event type name specified between the two tags of <T>. In this work, structural prompts are categorized into three groups: General Events, Overlapping Events, and Nested Events. The General Events describe the structural information of common events, guiding the PLM to identify potential interactions between frequently co-occurring events. The Overlapping Events describe the structural information of overlapping events, instructing the PLM to identify trigger words that commonly trigger overlapping events. The Nested Events describe the structural information of nested events, guiding the PLM to identify entities that frequently serve as argument roles in the events while acting as trigger words in others, or vice versa. Overall, the structural prompts are designed to describe the various structural information of biomedical events for generative event extraction.
@@ III-C Generative Model Training
The proposed GenBEE model employs a generative pre-trained language model BART [12] for encoding-decoding. In this module, the prefix $\mathcal{P}_{e_{i}}$ is integrated into both the encoder and decoder of the model. Specifically, from the prefix for event type $e_{i}$, we obtain the additional key and value matrices as $\mathcal{P}_{e_{i}}=K^{e_{i}}=V^{e_{i}}$, which can further be expressed as $K^{e_{i}}=\{k_{1}^{e_{i}},...,k_{l}^{e_{i}}\}$ and $V^{e_{i}}=\{v_{1}^{e_{i}},...,v_{l}^{e_{i}}\}$. Here, $k_{n}^{e_{i}}$ and $v_{n}^{e_{i}}$ $(n\in[1,l])$ are vectors with the same hidden dimension in the Transformer layer. These additional key and value matrices are concatenated with the original key and value matrices in the self-attention layers of the encoder and cross-attention layers of the decoder. As such, when computing the dot-product attention, the query matrix at each position is influenced by these structure-aware prefixes, which will subsequently influence the weightings of the representations for event generation. Note that the prefix $\mathcal{P}_{e_{i}}$ is learnable according to the querying event type $e_{i}$, thereby guiding the BART PLM to differentiate event types with different structural features.
@@
For training, the trainable parameters of the GenBEE model include those in the encoder-only pre-trained language model BioBERT and the generative pre-trained language model BART. The training objective of the GenBEE model is to minimize the negative log-likelihood of the ground truth sequence $\mathcal{Y^{\prime}}_{e_{i},j}$, given the output sequence $\mathcal{Y}_{e_{i},j}$:

### v2

# An Event Structure-aware Generative Model for Biomedical Event Extraction

@@ Abstract
Biomedical Event Extraction (BEE) is a challenging task that involves modeling complex relationships between fine-grained entities in biomedical text. Most existing BEE models rely on classification methods that ignore label semantics and argument dependencies in the data. Although generative models that use prompts are increasingly being used for event extraction, they face two main challenges: creating effective prompts for the biomedical domain and dealing with events with complex structures in the text. To address these limitations, we propose GenBEE, a generative model enhanced with structure-aware prefixes for biomedical event extraction. GenBEE constructs event prompts that leverage knowledge distilled from large language models (LLMs), thereby incorporating both label semantics and argument dependency relationships. Additionally, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes with structural prompts, enriching the generation process with structural features. Extensive experiments on three benchmark datasets demonstrate the effectiveness of GenBEE and it achieves state-of-the-art performance on the MLEE and GE11 datasets. Moreover, our analysis shows that the structural prefixes effectively bridge the gap between structural prompts and the representation space of generative models, enabling better integration of event structural information.
@@ I Introduction
Biomedical Event Extraction (BEE) is a challenging task that involves identifying molecular events from natural language text, where an event typically includes certain biomedical entities, such as genes, proteins, or cellular components [1]. BEE is typically divided into two subtasks, namely biomedical event trigger detection and biomedical event argument extraction. Biomedicalevent trigger detection aims to identify trigger words and classify them by event types that represent the presence of biomedical events. On the other hand, biomedical event argument extraction aims to identify the arguments and associate them with specific roles in biomedical events, correlating them with the corresponding event triggers. Moreover, BEE can provide invaluable information to facilitate the curation of knowledge bases and biomolecular pathways [2].
@@
Traditionally, BEE has been formulated as a classification problem [3, 4]. However, these classification-based methods often struggle with encoding label semantics and other forms of weak supervision information [5, 6, 7]. Moreover, many current methods extract events in a pipeline manner, where trigger detection and argument extraction are treated as separate phases, which hinders the learning of shared knowledge between subtasks [8]. Recently, generation-based models have been proposed for event extraction (EE) on general domains [9, 10]. These models typically cast the event extraction task as a sequence generation problem using encoder-decoder-based pre-trained language models (PLMs) [11, 12] to output conditional generation sequences. These generation-based models benefit from natural language prompts to learn additional semantic information [13]. Apart from employing natural language prompts to enhance semantic learning from the given context, some recent works such as GTEE [14] and AMPERE [15] have also proposed injecting continuous prompts [16, 17] into their frameworks for better contextual representation learning for event argument extraction.
@@
Figure 1: An example of nested events identified from the given text.
@@
However, most current generation-based EE models [14, 13, 15] have difficulties when dealing with events with complex structures, such as nested events. Compared to flat event structures, texts in the biomedical domain often contain more nested event structures [18]. For example, in benchmark datasets like MLEE [19] and GE11 [20], the nested event structures constitute approximately 25% and 28% of all events in the datasets, respectively. Figure 1 shows an example of two nested events identified from the given text. The trigger word “promoters” of the Positive Regulation event (E2) serves as the argument role Site for the trigger word “bind”, which triggers the Binding event (E1). Furthermore, “TCF-1 alpha” serves as the argument role Theme for both events (E1 and E2). In addition, a trigger word may also trigger different events in a sentence, depending on its event type or arguments. Events with such structures are called overlapping events. The rest of the events are regarded as flat events. Identifying these complex structures is non-trivial, as it can reveal the underlying relationships between events. For example, a Site2 argument in a Binding event is likely to be a trigger for a Positive Regulation event.
@@
For generation-based EE models, an intuitive way to incorporate structural information through prompt engineering [21] is by adding natural language prompts that describe complex structural information as part of the input to the generative PLM’s encoder. This helps elicit the rich knowledge distributed in the PLM. However, Fei et al. [22] pointed out that the mainstay contextual representations in a generative language model’s encoder tend to weaken other high-level information such as structural information, thereby reducing the effects of such information at the decoder. Moreover, it is also unclear whether generative PLMs, such as BART, contain sufficient domain knowledge to represent complex biomedical events. Therefore, incorporating structural information effectively into generative BEE models is challenging.
@@
To address the above issues, we propose an event structure-aware Generative Biomedical Event Extraction model (GenBEE), which can capture complex event structures in biomedical texts. First, GenBEE constructs type description prompts to integrate the semantics of type labels. Next, it distills knowledge from large language models (LLMs), such as GPT-4 [23] to create event template prompts, which incorporate information such as dependencies among argument roles, to enhance contextual representation learning. Then, GenBEE introduces a structural prefix learning module that generates structure-aware prefixes, enriched with structural prompts to infuse the structural features.
GenBEE functions as a comprehensive BEE framework that handles event trigger detection and event argument extraction simultaneously, and effectively leverages the shared knowledge and dependencies between these two subtasks. The main contributions of this paper are summarized as follows:
@@
  • We propose GenBEE, a structure-aware generative model for biomedical event extraction. Our proposed end-to-end model can distill knowledge from the LLM to incorporate label semantics, dependency relationships among arguments, and shared knowledge between subtasks.
@@
  • In the GenBEE model, we introduce a structural prefix learning module to guide the generation process. This module constructs the structure-aware prefix to link continuous prompts with the representation space of generative models. It is worth noting that this plug-and-play module can be added to any biomedical event extraction framework.
  • We have evaluated GenBEE on three widely-used BEE benchmark datasets, namely MLEE, GE11, and PHEE. Experimental results show that GenBEE has achieved state-of-the-art performance on the MLEE and GE11 datasets, and achieved competitive results when compared to the state-of-the-art classification-based models on the PHEE dataset.
@@ II-A Biomedical Event Extraction
Many current methods for biomedical event extraction adopt a “pipeline” strategy, where trigger detection and argument extraction are treated as separate phases. However, “pipeline” methods may result in the propagation of errors. To address this issue, Trieu et al. [8] introduced an end-to-end method called DeepEventMine, which simultaneously identifies triggers and allocates roles to entities, thereby mitigating error propagation. However, DeepEventMine requires the full annotation of all entities, which may not always be accessible in datasets. Ramponi et al. [6] considered biomedical event extraction as a sequence labeling task that involves joint training models for trigger detection and argument extraction via multi-task learning. Aside from these classification-based methods, some recent works have formulated event extraction as a question-answering task [24, 25]. This new paradigm transforms traditional classification methods into multiple questioning rounds, producing a natural language answer about a trigger or an argument in each round. Current QA-based event extraction methods are primarily focused on formulating separate questions for different events and argument types. However, these QA methods lack the capability on capturing dependency information between arguments, which could be helpful for assigning argument types.
@@ II-B Generative Event Extraction
Recently, another emerging line of work that treats event extraction as a conditional generation problem [26, 27, 28, 13] has evolved. The development of generation-based event extraction models begins with the reformulation of event extraction problems as a generation task. For example, TANL [26] treated event extraction as translation tasks with augmented labels using a variety of brackets and vertical bar symbols for representation, and tuned the model to predict these augmented labels. TempGen [27] proposed using templates for constructing role-filler tasks for entity extraction, and generating outputs that fill role entities into non-natural templated sequences. To adapt generation-based models to the task of event argument extraction, BART-Gen [28] investigated generative document-level event argument extraction. However, these methods only use naive natural language prompts and have not considered the incorporation of semantic information into the natural language prompts.
To better incorporate label semantics into the model, DEGREE [13] proposed a template-based model that incorporates label semantics into the designed prompts to extract event information. More recently, researchers have started to advance this line of work. For example, GTEE [14] used prefixes [16], which are sequences of tunable continuous prompts, to incorporate contextual type information into the generation-based model. AMPERE [15] proposed to equip the generative model with AMR-aware prefixes to embed the AMR graph, providing more contextual information. To the best of our knowledge, there is no work done on incorporating structural features of complex events into the generative models for BEE. Therefore, in this paper, we propose a novel generative model called GenBEE that integrates complex event structural information for BEE.
@@ III Method
In this section, we discuss our proposed GenBEE model in detail. The proposed GenBEE model performs generative biomedical event extraction as follows: The extraction process is divided into a number of subtasks based on each pre-defined event type of a dataset. Given a Dataset $\mathcal{D}$ $=$ $\left\{\mathcal{C}_{j}\mid j\in[1,|\mathcal{D}|]\right\}$ with its pre-defined event type set $\mathcal{E}$ $=$ $\left\{e_{i}\mid i\in[1,|\mathcal{E}|]\right\}$, where $\mathcal{C}_{j}$ denotes a textual context and $e_{i}$ denotes an event type. Each subtask is denoted as $S_{e_{i},\mathcal{C}}$, for the type of event $e_{i}$ and the context $\mathcal{C}$, where $\mathcal{C}$ contains all the textual contexts in the dataset. As such, we have a total of $|\mathcal{E}|$ generation subtasks based on each event type for all textual contexts $\mathcal{C}$. The proposed model conducts training iteratively based on each subtask with all textual contexts in the dataset. Figure 2 shows the architecture of the proposed model, which consists of the following modules: Event Prompt Construction, Structural Prefix Learning, and Generative Model Training.
@@ III-A Event Prompt Construction
Most previous works [14, 15] proposed to use manually designed templates for event extraction. However, it is challenging and time-consuming to define event templates for each event type from the biomedical event datasets, which requires domain knowledge. In this work, we propose to employ the Large Language Model (LLM) GPT-4 [23] with designed prompts to generate the event templates for the event types pre-defined in the biomedical event datasets. Figure 3 shows an input prompt that we have designed for generating the event template for the Binding event using GPT-4. The input prompt consists of an “Instruction” that specifies the generation objective for the biomedical event extraction template and a “Basic Template” that outlines the target trigger and argument roles, constructed based on the pre-defined roles given in the technical report associated with the biomedical event dataset. The output of GPT-4 is an updated event template for the corresponding event type, which contains richer information than the basic template for generative event extraction.
@@
where “;” denotes the sequence concatenation, and $[SEP]$ is the special separator token in the PLM (e.g., BART). For each subtask based on $e_{i}$, we construct the input tokens for all the contexts in the dataset.
@@ III-B Structural Prefix Learning
In biomedical texts, events are often complex and interrelated. Identifying the latent relationships among entities across different events could provide important clues for capturing complex event structures, thereby enhancing representation learning. Table II shows the structural prompts [29] constructed in the proposed model to capture different event structures for the Binding event. The structural prompts aim to provide descriptions of different event structures and then elicit the PLM (e.g., BioBERT) to identify complex event structures by capturing the latent relationships among entities across different biomedical events. As shown in Table II, the structural prompts are the same for all event types, except the corresponding event type name specified between the two tags of <T>. In this work, structural prompts are categorized into three groups: General Events, Overlapping Events, and Nested Events. The General Events describe the structural information of common events, guiding the PLM to identify potential interactions between frequently co-occurring events. The Overlapping Events describe the structural information of overlapping events, instructing the PLM to identify trigger words that commonly trigger overlapping events. The Nested Events describe the structural information of nested events, guiding the PLM to identify entities that frequently serve as argument roles in the events while acting as trigger words in others, or vice versa. Overall, the structural prompts are designed to describe the various structural information of biomedical events for generative event extraction.
@@ III-C Generative Model Training
The proposed GenBEE model employs a generative pre-trained language model BART [12] for encoding and decoding. In this module, the prefix $\mathcal{P}_{e_{i}}$ is integrated into both the encoder and decoder of the model. Specifically, from the prefix for event type $e_{i}$, we obtain the additional key and value matrices as $\mathcal{P}_{e_{i}}=K^{e_{i}}=V^{e_{i}}$, which can further be expressed as $K^{e_{i}}=\{k_{1}^{e_{i}},...,k_{l}^{e_{i}}\}$ and $V^{e_{i}}=\{v_{1}^{e_{i}},...,v_{l}^{e_{i}}\}$. Here, $k_{n}^{e_{i}}$ and $v_{n}^{e_{i}}$ $(n\in[1,l])$ are vectors with the same hidden dimension in the Transformer layer. These additional key and value matrices are concatenated with the original key and value matrices in the self-attention layers of the encoder and cross-attention layers of the decoder. As such, when computing the dot-product attention, the query matrix at each position is influenced by these structure-aware prefixes, which will subsequently influence the weightings of the representations for event generation. Note that the prefix $\mathcal{P}_{e_{i}}$ is learnable according to the querying event type $e_{i}$, thereby guiding the BART PLM to differentiate event types with different structural features.
@@
For training, the trainable parameters of the GenBEE model include those in the encoder-only pre-trained language model BioBERT and the generative pre-trained language model BART. The training objective of the GenBEE model is to minimize the negative log-likelihood of the ground truth sequence $\mathcal{Y^{\prime}}_{e_{i},j}$, given the output sequence $\mathcal{Y}_{e_{i},j}$ as follows: